# SReferences

### v1

@@ VIII Discussion
Markovian spin noise occurred during resource state generation is modelled by a depolarizing channel with spin error probability $p^{(s)}$ per photon generation time $\tau$. Our architecture can tolerate $p^{(s)}=0.15\%$, which translates to a spin coherence time of $T_{2}=500\tau$, already below the spin coherence times 333 With $\tau=5\times T_{\text{Mira}}\approx 68.9~{}$ns for quantum-dot entanglement sources [8] (determined by pulse-picking every $5$ picosecond pulses generated by the Ti:Sapphire Coherent Mira laser with repetition period $T_{\text{Mira}}=13.77$ ns), $T_{2}=34.5~{}\mu\text{s}<113~{}\mu$[40]. for some quantum emitters used in, i.e., spin-photon entanglement sources with an inherent spin-echo sequence [8].
Photon loss still remains to be the main bottleneck for FBQC, but recent experimental breakthroughs in photonic hardware [36, 37] and further optimization in the architecture hold promise to bridge this gap. For instance, coupling between different quantum emitters would enable deterministic entangling gates to significantly improve the logical fusion success probability and hence resistance to photon loss [15].
@@ Acknowledgement
I Erasure and error rates of Logical fusions
@@
Here we calculate the probability of obtaining or recovering logical outcomes $\overline{XX}$ and $\overline{ZZ}$ in a logical fusion, using REP and RUS in the presence of photon losses. The logical fusion error rates are also computed, with physical $X$ and $Z$ errors applied to each physical fusion. In general, for a physical fusion, with an end-to-end efficiency of $\eta$ and a fusion failure rate of $P_{\text{fail}}$ there are three possible events:
@@
The logical error rates of REP fusion can be found by counting the number of physical fusion errors and determining whether this gives a logical error. Let us assume that $X$ and $Z$ errors occur on each fusion outcome with probabilities $p_{X}$ and $p_{Z}$, which flip $ZZ$ and $XX$ outcomes, respectively. The $Z$ repetition code offers no assistance to the recovery of $\overline{XX}$ outcome, and so the probability of a logical error in $\overline{XX}$ is the probability of having an odd number of $Z$ errors in $m$ physical fusion attempts.
@@
Figure S1: Monte-Carlo simulation of RUS fusion. (a) Simulated logical fusion error as a function of end-to-end efficiency $\eta$. The simulation uses $N=11$ with $10^{4}$ trials. Analytical curves are obtained using Eq. (S5). (b) Distribution of number of fusion attempts $m$ in a RUS logical fusion. $z$-axis is the probability that the RUS fusion stops at $m$-th attempt. At $20\%$ loss, there is a $84\%$ chance that the RUS fusion terminates within $m=3$ attempts.
II RUS fusion with re-initialization of emitters

### v2

@@ VIII Discussion
Markovian spin noise occurred during resource state generation is modelled by a depolarizing channel with spin error probability $p^{(s)}$ per photon generation time $\tau$. Our architecture can tolerate $p^{(s)}=0.15\%$, which translates to a required spin coherence time of $T_{2}=500\tau$, already below the spin coherence times 222 With $\tau=5\times T_{\text{Mira}}\approx 68.9~{}$ns for quantum-dot entanglement sources [8] (determined by pulse-picking every $5$ picosecond pulses generated by the Ti:Sapphire Coherent Mira laser with repetition period $T_{\text{Mira}}=13.77$ ns), $T_{2}=34.5~{}\mu\text{s}<113~{}\mu$[40]. for some quantum emitters used in, i.e., spin-photon entanglement sources with an inherent spin-echo sequence [8].
Photon loss still remains to be the main bottleneck for FBQC, but recent experimental breakthroughs in photonic hardware [35, 36] and further optimization in the architecture hold promise to bridge this gap. For instance, deterministic entanglement between two emitters is possible by attempting RUS fusions between their respective coupled emitters, followed by entanglement swapping upon fusion success. This would significantly improve the logical fusion success probability and hence its resistance to photon loss [37, 15].
@@ Acknowledgement
I Proof for edge and node split rules
In this section, we provide proofs for the edge and node split rules using stabilizer formalism. We then show how the node split rule is applied to decompose the FFCC lattice in Fig. 1a into the synchronous FFCC fusion network in Fig. 1b.
Edge split rule is constructed based on the equivalence between a $\left\{XZ,ZX\right\}$ fusion and a $CZ$ gate between the two qubits involved in the fusion, before they are measured in the Pauli-$X$ basis. It can be applied to any two qubits connected with a common edge. In Fig. S1, we consider a general case where each of the qubits involved ($1$ and $2$) is connected to an arbitrary graph.
Figure S1: Edge split. $M_{X}$ is a $X$ measurement performed on the black qubit. Pink bar represents a $\left\{XZ,ZX\right\}$ fusion.
To show the equivalence of two graphs shown in Fig. S1, we compare their stabilizer generators after measurements. For LHS, the stabilizer generators associated with qubits $1$ and $2$ are:
$\mathcal{S}_{\text{LHS}}~{}=$ $~{}\langle\stackrel{{\scriptstyle*}}{{X_{1}Z_{\mathcal{N}_{1}\setminus\{2\}}Z_{2}}},~{}\{X_{j}Z_{\mathcal{N}_{j}\setminus\{1\}}Z_{1}\}_{j\in\mathcal{N}_{1}\setminus\{2\}},~{}X_{2}Z_{\mathcal{N}_{2}\setminus\{1\}}Z_{1},~{}\{X_{i}Z_{\mathcal{N}_{i}\setminus\{2\}}Z_{2}\}_{i\in\mathcal{N}_{2}\setminus\{1\}}\rangle$
$\xrightarrow{X_{2}}$$~{}\langle\{X_{j}Z_{\mathcal{N}_{j}\setminus\{1\}}Z_{1}\}_{j\in\mathcal{N}_{1}\setminus\{2\}},~{}\stackrel{{\scriptstyle*}}{{Z_{\mathcal{N}_{2}\setminus\{1\}}Z_{1}}},~{}\{X_{1}Z_{\mathcal{N}_{1}\setminus\{2\}}X_{i}Z_{\mathcal{N}_{i}\setminus\{2\}}\}_{i\in\mathcal{N}_{2}\setminus\{1\}}\rangle\otimes\langle X_{2}\rangle$
$\xrightarrow{X_{1}}$$~{}\langle\{Z_{\mathcal{N}_{2}\setminus\{1\}}X_{j}Z_{\mathcal{N}_{j}\setminus\{1\}}\}_{j\in\mathcal{N}_{1}\setminus\{2\}},~{}\{Z_{\mathcal{N}_{1}\setminus\{2\}}X_{i}Z_{\mathcal{N}_{i}\setminus\{2\}}\}_{i\in\mathcal{N}_{2}\setminus\{1\}}\rangle\otimes\langle X_{2}\rangle\otimes\langle X_{1}\rangle,$ (S1)
where $Z_{\mathcal{N}_{i}\setminus\{k\}}\equiv\prod_{j\in\mathcal{N}_{i}\setminus\{k\}}Z_{j}$ is the tensor product of Pauli-$Z$ operators on all neighboring qubits $\mathcal{N}_{i}$ connected to qubit $i$, but excluding the $Z$ operator on qubit $k$. The symbol ()(*)( ∗ ) indicates that the stabilizer generator anti-commutes with measurement in the subsequent step and is thus picked to construct a new list of post-measurement stabilizer generators [41]. For simplicity we assume that all measurements are deterministic and give positive outcomes. A negative outcome only changes the sign of stabilizers. The curly bracket $\{S_{j}\}_{j\in\mathcal{N}_{i}\setminus\{k\}}$ includes all stabilizer generators $S_{j}$ of qubit $j$ where $j$ belongs to a set of neighbors of qubit $i$ excluding qubit $k$.
For the graph on RHS, the relevant stabilizer generators evolve as
$\mathcal{S}_{\text{RHS}}~{}=$ $~{}\langle\stackrel{{\scriptstyle*}}{{X_{1^{\prime}}Z_{\mathcal{N}_{1^{\prime}}\setminus\{j\}}Z_{j}}},~{}\{X_{j}Z_{\mathcal{N}_{j}\setminus\{1^{\prime}\}}Z_{1^{\prime}}\}_{j\in\mathcal{N}_{1^{\prime}}},~{}X_{2^{\prime}}Z_{\mathcal{N}_{2^{\prime}}\setminus\{i\}}Z_{i},~{}\{X_{i}Z_{\mathcal{N}_{i}\setminus\{2^{\prime}\}}Z_{2^{\prime}}\}_{i\in\mathcal{N}_{2^{\prime}}}\rangle$
$\xrightarrow{Z_{1^{\prime}}X_{2^{\prime}}}$$~{}\langle\{X_{j}Z_{\mathcal{N}_{j}\setminus\{1^{\prime}\}}Z_{1^{\prime}}\}_{j\in\mathcal{N}_{1^{\prime}}},~{}\stackrel{{\scriptstyle*}}{{X_{2^{\prime}}Z_{\mathcal{N}_{2^{\prime}}\setminus\{i\}}Z_{i}}},~{}\{X_{1^{\prime}}Z_{2^{\prime}}Z_{\mathcal{N}_{1^{\prime}}\setminus\{j\}}Z_{j}X_{i}Z_{\mathcal{N}_{i}\setminus\{2^{\prime}\}}\}_{i\in\mathcal{N}_{2^{\prime}}}\rangle\otimes\langle Z_{1^{\prime}}X_{2^{\prime}}\rangle$
$\xrightarrow{X_{1^{\prime}}Z_{2^{\prime}}}$$~{}\langle\{Z_{\mathcal{N}_{2^{\prime}}\setminus\{i\}}Z_{i}X_{j}Z_{\mathcal{N}_{j}\setminus\{1^{\prime}\}}\}_{j\in\mathcal{N}_{1^{\prime}}},~{}\{Z_{\mathcal{N}_{1^{\prime}}\setminus\{j\}}Z_{j}X_{i}Z_{\mathcal{N}_{i}\setminus\{2^{\prime}\}}\}_{i\in\mathcal{N}_{2^{\prime}}}\rangle\otimes\langle Z_{1^{\prime}}X_{2^{\prime}}\rangle\otimes\langle X_{1^{\prime}}Z_{2^{\prime}}\rangle$
$=$ $~{}\langle\{Z_{\mathcal{N}_{2}\setminus\{1\}}X_{j}Z_{\mathcal{N}_{j}\setminus\{1\}}\}_{j\in\mathcal{N}_{1}\setminus\{2\}},~{}\{Z_{\mathcal{N}_{1}\setminus\{2\}}X_{i}Z_{\mathcal{N}_{i}\setminus\{2\}}\}_{i\in\mathcal{N}_{2}\setminus\{1\}}\rangle\otimes\langle Z_{1^{\prime}}X_{2^{\prime}}\rangle\otimes\langle X_{1^{\prime}}Z_{2^{\prime}}\rangle,$ (S2)
which gives identical stabilizer generators as in Eq. (S1). The equality in the last step results from the fact that the set of neighboring qubits $\mathcal{N}_{1^{\prime}}$ ($\mathcal{N}_{2^{\prime}}$) is equivalent to the set $\mathcal{N}_{1}\setminus\{2\}$ ($\mathcal{N}_{2}\setminus\{1\}$). $X_{1^{\prime}}Z_{2^{\prime}}$ and $Z_{1^{\prime}}X_{2^{\prime}}$ are measurement operators for a successful $\{XZ,ZX\}$ physical fusion on qubits $1^{\prime}$ and $2^{\prime}$.
Node split rule states that measuring a qubit of a cluster state in the Pauli-$X$ basis is equivalent to applying a $\left\{XX,ZZ\right\}$ fusion (yellow bar in Fig. S2) between two additional qubits introduced to the same node of the graph.
Figure S2: Node split. Qubits $i$ and $j$ are included to show equivalence in the surviving stabilizer group between both graphs.
The graph on LHS is stabilized by generators which change after measuring qubit $1$ in the $X$ basis:
$\mathcal{S}_{\text{LHS}}=$ $~{}\langle X_{1}Z_{\mathcal{N}_{1}},~{}\stackrel{{\scriptstyle*}}{{X_{j}Z_{\mathcal{N}_{j}\setminus\{1\}}Z_{1}}},~{}X_{i}Z_{\mathcal{N}_{i}\setminus\{1\}}Z_{1},~{}\{X_{k}Z_{\mathcal{N}_{k}}Z_{1}\}_{k\in\mathcal{N}_{1}\setminus\{j,i\}}\rangle$
$\xrightarrow{X_{1}}$ $~{}\langle Z_{\mathcal{N}_{1}},~{}X_{j}Z_{\mathcal{N}_{j}\setminus\{1\}}X_{i}Z_{\mathcal{N}_{i}\setminus\{1\}},~{}\{X_{j}Z_{\mathcal{N}_{j}\setminus\{1\}}X_{k}Z_{\mathcal{N}_{k}}\}_{k\in\mathcal{N}_{1}\setminus\{j,i\}}\rangle\otimes\langle X_{1}\rangle.$ (S3)
Similarly, for RHS,
$\mathcal{S}_{\text{RHS}}=$ $~{}\langle\stackrel{{\scriptstyle*}}{{X_{1_{a}}Z_{\mathcal{N}_{1_{a}}\setminus\{j\}}Z_{j}}},~{}X_{j}Z_{\mathcal{N}_{j}\setminus\{1_{a}\}}Z_{1_{a}},~{}\{X_{k}Z_{\mathcal{N}_{k}\setminus\{1_{a}\}}Z_{1_{a}}\}_{k\in\mathcal{N}_{1_{a}}\setminus\{j\}},\quad$
$\quad\quad~{}X_{1_{b}}Z_{\mathcal{N}_{1_{b}}\setminus\{i\}}Z_{i},~{}X_{i}Z_{\mathcal{N}_{i}\setminus\{1_{b}\}}Z_{1_{b}},~{}\{X_{k}Z_{\mathcal{N}_{k}\setminus\{1_{b}\}}Z_{1_{b}}\}_{k\in\mathcal{N}_{1_{b}}\setminus\{i\}}\rangle$
$\xrightarrow{Z_{1_{a}}Z_{1_{b}}}$ $~{}\langle\stackrel{{\scriptstyle*}}{{X_{j}Z_{\mathcal{N}_{j}\setminus\{1_{a}\}}Z_{1_{a}}}},~{}\{X_{k}Z_{\mathcal{N}_{k}\setminus\{1_{a}\}}Z_{1_{a}}\}_{k\in\mathcal{N}_{1_{a}}\setminus\{j\}},~{}X_{1_{a}}X_{1_{b}}Z_{\mathcal{N}_{1_{a}}\setminus\{j\}}Z_{j}Z_{\mathcal{N}_{1_{b}}\setminus\{i\}}Z_{i},$
$\quad\quad~{}X_{i}Z_{\mathcal{N}_{i}\setminus\{1_{b}\}}Z_{1_{b}},~{}\{X_{k}Z_{\mathcal{N}_{k}\setminus\{1_{b}\}}Z_{1_{b}}\}_{k\in\mathcal{N}_{1_{b}}\setminus\{i\}}\rangle\otimes\langle Z_{1_{a}}Z_{1_{b}}\rangle$
$\xrightarrow{X_{1_{a}}X_{1_{b}}}$$~{}\langle\{X_{j}Z_{\mathcal{N}_{j}\setminus\{1_{a}\}}X_{k}Z_{\mathcal{N}_{k}\setminus\{1_{a}\}}\}_{k\in\mathcal{N}_{1_{a}}\setminus\{j\}},~{}Z_{\mathcal{N}_{1_{a}}\setminus\{j\}}Z_{j}Z_{\mathcal{N}_{1_{b}}\setminus\{i\}}Z_{i},$
$\quad\quad~{}X_{j}Z_{\mathcal{N}_{j}\setminus\{1_{a}\}}X_{i}Z_{\mathcal{N}_{i}\setminus\{1_{b}\}},~{}\{X_{j}Z_{\mathcal{N}_{j}\setminus\{1_{a}\}}X_{k}Z_{\mathcal{N}_{k}\setminus\{1_{b}\}}\}_{k\in\mathcal{N}_{1_{b}}\setminus\{i\}}\rangle\otimes\langle Z_{1_{a}}Z_{1_{b}}\rangle\otimes\langle X_{1_{a}}X_{1_{b}}\rangle$
$=$ $~{}\langle Z_{\mathcal{N}_{1}},~{}X_{j}Z_{\mathcal{N}_{j}\setminus\{1\}}X_{i}Z_{\mathcal{N}_{i}\setminus\{1\}},~{}\{X_{j}Z_{\mathcal{N}_{j}\setminus\{1\}}X_{k}Z_{\mathcal{N}_{k}}\}_{k\in\mathcal{N}_{1}\setminus\{j,i\}}\rangle\otimes\langle Z_{1_{a}}Z_{1_{b}}\rangle\otimes\langle X_{1_{a}}X_{1_{b}}\rangle.$ (S4)
The equality in the last step originates from (1) $Z_{\mathcal{N}_{1}}=Z_{\mathcal{N}_{1_{a}}}Z_{\mathcal{N}_{1_{b}}}$, and (2) the set of qubits in $\mathcal{N}_{j}\setminus\{1\}$ ($\mathcal{N}_{i}\setminus\{1\}$) of LHS is identical to the set $\mathcal{N}_{j}\setminus\{1_{a}\}$ ($\mathcal{N}_{i}\setminus\{1_{b}\}$).
Decomposition into synchronous FFCC fusion network using node splits and $X$ measurements
The decomposition can be understood in two steps: First, apply node split to every check qubit (white circle) of the lattice in Fig. 1a. Next, $X$ measurements are performed on two neighboring qubits previously connected to the check qubit.
In Fig. S3a we visualize the decomposition process in the graph state picture. We begin with a ladder-like graph in the FFCC lattice. As described in the main text, in the measurement-based model, all qubits including data (black) and check qubits (white) are measured in the Pauli-$X$ basis to propagate information from bottom layers to the top. Since we are interested in the graph transformation in the vicinity of the check qubit $i$, only $X$ measurements on its neighbors (qubits $2$ and $6$) which are relevant to the transformation are shown in the first graph of Fig. S3. The idea of decomposition is to split the ladder-like graph into a fusion between two simpler states, i.e., linear cluster states. Therefore, we choose to split qubit $i$ into two additional qubits, instead of its neighbors. Using the node split rule in Fig. S2, we obtain the second graph with additional qubits $4$ and $8$, which is equivalent to a fusion between two branched chain states followed by $X$ measurements on qubits $2$ and $6$.
Figure S3: Graph decomposition using node split and $X$ measurements. Here we consider the decomposition of a section of the FFCC lattice into a fusion between linear cluster states. First, the check qubit (white) is split into qubits $4$ and $8$ which will undergo a $\{XX,ZZ\}$ fusion. Second, $X$ measurements on qubits $2$ and $6$ remove them from the graph and connect their neighbors, with a Hadamard gate on each of the neighbors.
Next, instead of applying node split again, these $X$ measurements are performed to transform the graph into two linear cluster states, with Hadamard gates on both qubits $4$ and $8$, as shown in the third graph of Fig. S3. These Hadamard gates do not change the fusion measurement outcomes and only alter the fusion failure basis. The graph transformation under a $X$ measurement can be verified similarly. After measuring qubit $2$ in $X$ in the second graph of Fig. S3, the stabilizer generators of the left branch chain state become
$\mathcal{S}=$ $~{}\langle X_{1}Z_{2}Z_{\mathcal{N}_{1}\setminus\{2\}},~{}X_{2}Z_{1}Z_{3}Z_{4},~{}X_{3}Z_{2}Z_{\mathcal{N}_{3}\setminus\{2\}},~{}\stackrel{{\scriptstyle*}}{{X_{4}Z_{2}}}\rangle$
$\xrightarrow{X_{2}}$ $~{}\langle X_{1}X_{4}Z_{\mathcal{N}_{1}\setminus\{2\}},~{}Z_{1}Z_{3}Z_{4},~{}X_{3}X_{4}Z_{\mathcal{N}_{3}\setminus\{2\}}\rangle$
$=$ $~{}H_{4}\bigg{[}\langle X_{1}Z_{4}Z_{\mathcal{N}_{1}\setminus\{2\}},~{}Z_{1}Z_{3}X_{4},~{}X_{3}Z_{4}Z_{\mathcal{N}_{3}\setminus\{2\}}\rangle\bigg{]},$ (S5)
which is indeed identical to a linear cluster state with a Hadamard rotation on qubit $4$.
Therefore, repeating the above steps on all other check qubits in Fig. 1a, we obtain the synchronous FFCC fusion network depicted in Fig. 1b. It is important to emphasize that during resource state generation, these $X$ measurements are not implemented, i.e., they are effectively virtual measurements that always give $+1$ outcome. As a result of the equivalence depicted in Fig. S3, it is only necessary to generate linear cluster states and perform fusion measurements. In other words, the synchronous FFCC fusion network is equivalent to the FFCC lattice in the measurement-based model, where these $X$ measurements are always successful resulting in the $+1$ outcome.
As a final check, the equivalence between the first and third graphs in Fig. S3 is verified by comparing their surviving stabilizer groups:
$\mathcal{S}_{\text{LHS}}=$ $~{}\langle X_{1}Z_{\mathcal{N}_{1}\setminus\{2\}}Z_{2},~{}\stackrel{{\scriptstyle*}}{{X_{2}Z_{1}Z_{i}Z_{3}}},~{}X_{3}Z_{\mathcal{N}_{3}\setminus\{2\}}Z_{2},~{}X_{i}Z_{2}Z_{6},~{}X_{6}Z_{5}Z_{i}Z_{7},~{}X_{5}Z_{\mathcal{N}_{5}\setminus\{6\}}Z_{6},~{}X_{7}Z_{\mathcal{N}_{7}\setminus\{6\}}Z_{6}\rangle$
$\xrightarrow{X_{i}}$ $~{}\langle\stackrel{{\scriptstyle*}}{{X_{1}Z_{\mathcal{N}_{1}\setminus\{2\}}Z_{2}}},~{}X_{3}Z_{\mathcal{N}_{3}\setminus\{2\}}Z_{2},~{}Z_{2}Z_{6},~{}X_{2}Z_{1}Z_{3}X_{6}Z_{5}Z_{7},~{}X_{5}Z_{\mathcal{N}_{5}\setminus\{6\}}Z_{6},~{}X_{7}Z_{\mathcal{N}_{7}\setminus\{6\}}Z_{6}\rangle\otimes\langle X_{i}\rangle$
$\xrightarrow{X_{2}}$ $~{}\langle X_{1}Z_{\mathcal{N}_{1}\setminus\{2\}}X_{3}Z_{\mathcal{N}_{3}\setminus\{2\}},~{}\stackrel{{\scriptstyle*}}{{X_{1}Z_{\mathcal{N}_{1}\setminus\{2\}}Z_{6}}},~{}Z_{1}Z_{3}X_{6}Z_{5}Z_{7},~{}X_{5}Z_{\mathcal{N}_{5}\setminus\{6\}}Z_{6},~{}X_{7}Z_{\mathcal{N}_{7}\setminus\{6\}}Z_{6}\rangle\otimes\langle X_{i}\rangle\otimes\langle X_{2}\rangle$
$\xrightarrow{X_{6}}$ $~{}\langle X_{1}Z_{\mathcal{N}_{1}\setminus\{2\}}X_{3}Z_{\mathcal{N}_{3}\setminus\{2\}},~{}Z_{1}Z_{3}Z_{5}Z_{7},~{}X_{1}Z_{\mathcal{N}_{1}\setminus\{2\}}X_{5}Z_{\mathcal{N}_{5}\setminus\{6\}},~{}X_{1}Z_{\mathcal{N}_{1}\setminus\{2\}}X_{7}Z_{\mathcal{N}_{7}\setminus\{6\}}\rangle\otimes\langle X_{i}\rangle\otimes\langle X_{2}\rangle\otimes\langle X_{6}\rangle.$ (S6)
$\mathcal{S}_{\text{RHS}}=$ $~{}\langle\stackrel{{\scriptstyle*}}{{X_{1}Z_{\mathcal{N}_{1}\setminus\{4\}}Z_{4}}},~{}X_{4}Z_{1}Z_{3},~{}X_{3}Z_{\mathcal{N}_{3}\setminus\{4\}}Z_{4},~{}X_{5}Z_{\mathcal{N}_{5}\setminus\{8\}}Z_{8},~{}X_{8}Z_{5}Z_{7},~{}X_{7}Z_{\mathcal{N}_{7}\setminus\{8\}}Z_{8}\rangle$
$\xrightarrow{X_{4}X_{8}}$ $~{}\langle\stackrel{{\scriptstyle*}}{{X_{4}Z_{1}Z_{3}}},~{}X_{1}Z_{\mathcal{N}_{1}\setminus\{4\}}X_{3}Z_{\mathcal{N}_{3}\setminus\{4\}},~{}X_{1}Z_{\mathcal{N}_{1}\setminus\{4\}}X_{5}Z_{\mathcal{N}_{5}\setminus\{8\}}Z_{4}Z_{8},~{}X_{8}Z_{5}Z_{7},~{}X_{1}Z_{\mathcal{N}_{1}\setminus\{4\}}X_{7}Z_{\mathcal{N}_{7}\setminus\{8\}}Z_{4}Z_{8}\rangle\otimes\langle X_{4}X_{8}\rangle$
$\xrightarrow{Z_{4}Z_{8}}$ $~{}\langle X_{1}Z_{\mathcal{N}_{1}\setminus\{4\}}X_{3}Z_{\mathcal{N}_{3}\setminus\{4\}},~{}X_{1}Z_{\mathcal{N}_{1}\setminus\{4\}}X_{5}Z_{\mathcal{N}_{5}\setminus\{8\}},~{}Z_{1}Z_{3}Z_{5}Z_{7},~{}X_{1}Z_{\mathcal{N}_{1}\setminus\{4\}}X_{7}Z_{\mathcal{N}_{7}\setminus\{8\}}\rangle\otimes\langle X_{4}X_{8}\rangle\otimes\langle Z_{4}Z_{8}\rangle,$ (S7)
with $Z_{\mathcal{N}_{1}\setminus\{2\}}$ ($Z_{\mathcal{N}_{7}\setminus\{6\}}$) in the first graph is identical to $Z_{\mathcal{N}_{1}\setminus\{4\}}$ ($Z_{\mathcal{N}_{7}\setminus\{8\}}$) in the final graph.
II Erasure and error rates of Logical fusions
@@
Here we calculate the probability of obtaining or recovering logical outcomes $\overline{XX}$ and $\overline{ZZ}$ in a logical fusion, using REP and RUS in the presence of photon losses. The logical fusion error rates are also computed, with $X$ and $Z$ errors applied to each physical fusion. In general, with an end-to-end efficiency of $\eta$ and a fusion failure rate of $P_{\text{fail}}$, for a physical fusion there are three possible events:
@@
The logical error rates of REP fusion can be found by counting the number of physical fusion errors and determining whether this gives a logical error. Let us assume that $X$ and $Z$ errors occur on each fusion outcome with probabilities $p_{X}$ and $p_{Z}$, which flip $ZZ$ and $XX$ outcomes, respectively. If the probability that a photon acquiring $Z$ error prior to fusion is $e_{Z}$, then the probability of having a flipped $XX$ fusion outcome is the probability of odd numbers of photon $Z$ error which is $p_{X}=2e_{Z}$. Similarly on the logical level, since the $Z$ repetition code offers no assistance to the recovery of $\overline{XX}$ outcome, the probability of a logical error in $\overline{XX}$ is the probability of having an odd number of $Z$ fusion errors in $m$ physical fusion attempts:
@@
Figure S4: Monte-Carlo simulation of RUS fusion. (a) Simulated logical fusion success probability as a function of end-to-end efficiency $\eta$. The simulation uses $N=11$ with $10^{4}$ trials. Analytical curves are obtained using Eq. (S12). (b) Distribution of number of fusion attempts $m$ in a RUS logical fusion. $z$-axis is the probability that the RUS fusion stops at $m$-th attempt. At $20\%$ loss, there is a $84\%$ chance that the RUS fusion terminates within $m=3$ attempts.
III RUS fusion with re-initialization of emitters